A 7.0 * 10-9 C point charges at the center of a 5.0 cm radius sphear. What is the electric flux through the sphere surface?


1
Expert's answer
2021-03-01T12:39:32-0500

Electric field of sphere = "\dfrac{Q}{4\pi\epsilon_{\omicron}R^2}"

Surface area of sphere = "4\pi R^2"

Flux "=" "E\times area"

"\space\space\space\space\space\space\space=\dfrac{Q}{4\pi\epsilon_{\omicron}R^2}\times4\pi R^2"

"\space\space\space\space\space\space\space=\dfrac{Q}{\epsilon_{\omicron}}"

"\space\space\space\space\space\space\space=\dfrac{7\times10^{-9}}{8.8\times10^{-12}}=7.95\times10^2\space Nm^2/C"